### v2

# Topological materials with extensive flat-band surface states

Abstract
Materials that have zero-energy flat band states on the surface are likely to show surface superconductivity. Here we report a theoretical observation that a Hamiltonian describing a thin slab of topological nodal line semimetal, has zero energy eigenstate spanning the entire surface under certain conditions, namely (i) the hopping probability of fermions in the direction of thickness is more than that in other directions (ii) the onsite interaction should be less than some limiting value determined by the hopping probability. Our claim is substantiated by analytic and numerical approach. We also report new phase transitions in a region of parameter space and indicate that the Hamiltonian can also be realised by stacked layers described by Kitaev model on honeycomb lattice.
Topological nodal line semimetal (TNLSM) is a type of topological material characterised by the presence of nodal points (where the conduction and valence bands touch) along a closed loop in the momentum space. These materials have a unique importance due to their distinctive properties, like flat drumhead states [1, 2, 3], graphene-like transport properties in 3D [4] and plasmon mode behaviour [5, 6]. Bulk superconductivity has been observed in several such materials, such as, CaSb2 [7], SnTaS2 [8], NaAlSi [9], PbTaSe2 [10] . Surface superconductivity in a topological flat band system has been predicted long back [11, 12, 13, 14, 15, 16], and has been observed for graphene multilayers [17] and the TNLSM system Pd-doped CaAgP [18]. Surface superconductivity has also been observed for 3D Dirac material BaMg2Bi2 [19] and Weyl semimetal PtBi2 [20]. Furthermore, it has been shown by ab initio calculations [21] that the topological materials Y2C and Sc2C should have zero energy states covering extended regions of Brilloun zone surface.
In this article, we report a reasonable Hamiltonian which shows flat band surface states on the top and bottom surfaces of a sample with thin slab geometry, encompassing the entire X-Y plane in momentum space. Thus, a material prepared to be represented by this Hamiltonian will have the prospect of having prominent surface superconductivity with a high critical temperature. This model shows a rich phase diagram consisting of a number of topological phases. The Hamiltonian under question in the 3D wave-vector space (with periodic boundary condition) is constituted of $2\times 2$ free spinless Fermions:
${\mathcal{H}}=\sum_{\vec{k}}c_{\vec{k}}^{\dagger}{\mathcal{H}}_{\vec{k}}c_{\vec{k}}\;\;\;\;\;\;{\mathcal{H}}_{\vec{k}}=A_{\vec{k}}\sigma_{z}+B_{\vec{k}}\sigma_{x}$ (1)
with
$A_{\vec{k}}=J-t\cos k_{x}-t\cos k_{y}-\cos k_{z}$ (2)
$B_{\vec{k}}=t\sin k_{x}+t\sin k_{y}+\sin k_{z}$
where, $c_{\vec{k}}^{\dagger}=(c_{\vec{k}A}^{\dagger}\;\;\;c_{\vec{k}B}^{\dagger})$, $c_{\vec{k}}=(c_{\vec{k}A}\;\;\;c_{\vec{k}B})$ denote the creation and annihilation operators for two sublattices $A$ and $B$, $J$ and $t$ are parameters and $\vec{k}$ runs over the first Brillouin zone (BZ) of a cubic lattice $-\pi<k_{x},k_{y},k_{z}\leq\pi$. (The motivations for this Hamiltonian will be discussed below.) Note that this Hamiltonian is anisotropic in the Z direction (unless $t=1$)
Figure 1: The phase diagram (at zero temperature) in $t-J$ plane. There are six phases, irrespective of whether we have open or closed boundary in the Z direction. Fig. 1a is for periodic boundary in X and Y direction and open boundary in Z direction. We show here the regions on the toroidal $k_{x}-k_{y}$ surface at which zero-energy states appear at $n_{z}=\pm L$. Fig. 1b is for fully periodic boundaries. We show here the nodal lines (where the energy is zero) in the $k_{x}-k_{y}-k_{z}$ Brillouin zone. In both the figures, the nature of non-analyticity at the phase boundary is indicated by colour, dashed magenta corresponds to divergence of $\partial^{3}E_{0}/\partial t^{3}$ and red, to a discontinuous change in the same quantity.For computing the location of nodal lines we used $N=400^{3}$. For computing the surface states, we used $L_{z}=200$. Phase I: Topological insulator with zero-energy surface state ranging all over the $k_{x}-k_{y}$ plane. Phase II: Nodal line semimetal phase, with one nodal line, which touches $k_{x}-k_{z}$ and $k_{y}-k_{z}$ boundary surfaces. The zero-energy states form a multiply connected region, but the non-zero energy states form a simply connected region (a patch). Phase III: Another nodal line phase, but with two nodal lines, which touch all the surfaces. Both the zero-energy and non-zero-energy states form multiply connected regions. Phase IV: Two nodal lines in this phase touch only the $k_{x}-k_{y}$ boundary surfaces, but not the $k_{x}-k_{z}$ and $k_{y}-k_{z}$ surfaces. The non-zero-energy states form a multiply connected region, but the zero-energy states form two patches. Phase V: This phase contains only one nodal loop, which does not touch any boundary surface of the BZ. The non-zero-energy states form a multiply connected region, but the zero-energy states form one patch. Phase VI: This is a normal insulator phase without any node and zero-energy state on surface.
The corresponding Hamiltonian in real space (with periodic boundary condition) is (Eq. 3),
$\mathcal{H}=\sum_{n_{x}=-L_{x}}^{L_{x}}\sum_{n_{y}=-L_{y}}^{L_{y}}\sum_{n_{z}=-L_{z}}^{L_{z}}\left[c_{n_{x}n_{y}n_{z}}^{\dagger}J\sigma_{z}c_{n_{x}n_{y}n_{z}}+tc_{n_{x}n_{y}n_{z}}^{\dagger}\sigma^{\pm}c_{n_{x}n_{y}\pm 1\,n_{z}}+tc_{n_{x}n_{y}n_{z}}^{\dagger}\sigma^{\pm}c_{n_{x}\pm 1\,n_{y}n_{z}}\right.$
$\left.+c_{n_{x}n_{y}n_{z}}^{\dagger}\sigma^{\pm}c_{n_{x}n_{y}n_{z}\pm 1}\right]$ (3)
where $\sigma^{\pm}=-(\sigma_{z}\pm i\sigma_{x})/2$. This Hamiltonian represents a system with on-site energy $J$ and hops to and from the nearest neighbouring sites along X, Y, Z axes with amplitudes $t$, $t$, 1 respectively. To adapt the forms of Eqs (1,3) to the slab geometry, we perform inverse Fourier transform of the BZ Hamiltonian Eq. (1) in the Z direction and set open boundary condition in that direction. An alternative is to perform Fourier transformation of the direct space Hamiltonian Eq. (3) in X and Y directions and impose open boundary condition along Z direction. This gives (Eq. 4),
$\mathcal{H}=\sum_{k_{x},k_{y}}\sum_{n_{z}=-L_{z}}^{L_{z}}\left[c_{k_{x}k_{y}n_{z}}^{\dagger}\left(\Delta_{1}\sigma_{z}+\Delta_{2}\sigma_{x}\right)c_{k_{x}k_{y}n_{z}}+c_{k_{x}k_{y}n_{z}}^{\dagger}\sigma^{+}c_{k_{x}k_{y}n_{z}+1}+c_{k_{x}k_{y}n_{z}}^{\dagger}\sigma^{-}c_{k_{x}k_{y}n_{z}-1}\right]$ (4)
where $\Delta_{1}=J-t\cos(k_{x})-t\cos(k_{y})$, and $\Delta_{2}=t\sin(k_{x})+t\sin(k_{y})$. (For each $n_{z}$, the $k_{x}-k_{y}$ surface is then a toroid.) A similar Hamiltonian has been proposed for layered materials by Burkov, Hook and Balents [4]. The crucial result of this paper is that the zero energy eigenstates of this Hamiltonian (Eq (4)),which are localized on the surfaces $n_{z}=1$ and $n_{z}=L$, cover the entire $k_{x}-k_{y}$ face of the BZ for the parameter region $10<J<$ 1, $20<t<(1-J)/$ 2 and form a flat band. Hence a material described by our Hamiltonian is likely to show high-temperature surface superconductivity, as predicted elsewhere [11]. Furthermore, the existence of flat bands all-over the surface, is not destroyed if the hopping amplitude in Z direction fluctuates around 1 (and is uniformly distributed between 0.9 and 1.1).
Motivations for the Hamiltonian: We have mentioned above that our Hamiltonian (3) has been used to describe layered materials. For the special case of $t=1$, there are several other motivations:
(i) On the plane $k_{x}=-k_{y}$, for small $\vec{k}$, our Hamiltonian of Eq. (1) reduces to the usual Hamiltonian for TNLSM upto terms quadratic in $\vec{k}$ [22], [23], [24].
${\mathcal{H}}_{\vec{k}}^{0}=vk_{z}\sigma_{x}+(k^{2}-k_{0}^{2})\sigma_{z}$ (5)
(Here $v$ and $k_{0}$ are the parameters of the model and insulator to semimetal transition occurs as the parameter $k_{0}$ is varied.)
(ii) The usual Hamiltonian for Weyl semimetal [22, 23, 25, 26]
${\mathcal{H}}_{\vec{k}}^{W}=(J-\cos k_{x}-\cos k_{y}-\cos k_{z})\sigma_{z}$ (6)
$+\sin k_{x}\sigma_{x}+\sin k_{y}\sigma_{y}$
has a form similar to Eq. (1).
(iii) Kitaev model on honeycomb lattice can be represented [25, 26] by Eq. (1) with
$A_{\vec{k}}=J-\cos k_{x}-\cos k_{y}\;\;\;\;B_{\vec{k}}=\sin k_{x}+\sin k_{y}$ (7)
and a straightforward generalisation of these equations lead to Eq. (Topological materials with extensive flat-band surface states).
(iv) The matrix representation for the Hamiltonian (4) has a tridiagonal form
$\left(\begin{array}[]{cccccc}S&\sigma^{-}&0&0&\cdots&0\\ \sigma^{+}&S&\sigma^{-}&0&\cdots&0\\ 0&\sigma^{+}&S&\sigma^{-}&\cdots&0\\ \cdots&&&&&\\ 0&0&0&\cdots&\sigma^{+}&S\end{array}\right)$ (8)
where $S=\Delta_{1}\sigma_{z}+\Delta_{2}\sigma_{x}$. Since $S$ corresponds to the Kitaev Hamiltonian in Eq. (7), our Hamiltonian can also be realised by stacking layers of Kitaev honeycomb layers with intralayer hopping probability $t$ and interlayer hopping probability unity.
Symmetries of the Hamiltonian: The time reversal operator is ${\mathcal{T}}=\mathbf{1}{\mathcal{K}}$ (where ${\mathcal{K}}$ is complex conjugation operator [27]) and we observe that,
${\mathcal{T}}{\mathcal{H}}_{\vec{k}}{\mathcal{T}}^{-1}={\mathcal{H}}_{-\vec{k}}=A_{-\vec{k}}\sigma_{z}+B_{-\vec{k}}\sigma_{x}$ (9)
$=A_{\vec{k}}\sigma_{z}-B_{\vec{k}}\sigma_{x}$
${\mathcal{T}}(A_{\vec{k}}\sigma_{z}+B_{\vec{k}}\sigma_{x}){\mathcal{T}}^{-1}=A_{\vec{k}}\sigma_{z}+B_{\vec{k}}\sigma_{x}$
Since the Brillouin zone is symmetric about $\vec{k}=0$, corresponding to every $\vec{k}$ in the sum of Eq. (1) there will be a $-\vec{k}$. Hence our Hamiltonian is time-reversal symmetric. By the same argument, the Hamiltonian is also invariant under inversion symmetry $\mathcal{P}=\sigma_{z}$ since, ${\sigma_{z}}\mathcal{H}(-\vec{k}){\sigma_{z}}=\mathcal{H}(\vec{k})$. There is sub-lattice symmetry also, since $\sigma_{y}\mathcal{H}(\vec{k})\sigma_{y}=-\mathcal{H}(\vec{k})$.
The $\mathcal{H}_{\vec{k}}$ of the model under study is a real Hamiltonian, which is a consequence of the presence of both time reversal and inversion symmetry. The presence of $\mathcal{P}$ and $\mathcal{T}$ symmetries implies that the model is ${\mathcal{P}}*{\mathcal{T}}$ symmetric. It can also be proved that $({\mathcal{P}}*{\mathcal{T}})^{2}=1$ and ${\mathcal{P}}*{\mathcal{T}}$ is an anti-unitary operator, which in turn implies that ${\mathcal{P}}*{\mathcal{T}}$ can always be represented as $\mathcal{K}$ in a suitable basis. In that basis, the Hamiltonian will always be real.
Nodal Lines and Phase Diagram: These lines consist of the points at which energy is zero, that is, $A_{\vec{k}}=0$ and $B_{\vec{k}}=0$, with $A_{\vec{k}}$, $B_{\vec{k}}$ given by Eq. (Topological materials with extensive flat-band surface states). For a given $k_{z}$, one obtains,
$2t\cos\frac{k_{x}+k_{y}}{2}\cos\frac{k_{x}-k_{y}}{2}=J-\cos k_{z}$ (10)
$2t\sin\frac{k_{x}+k_{y}}{2}\cos\frac{k_{x}-k_{y}}{2}=-\sin k_{z}.$
and one can solve out $k_{x}$, $k_{y}$ from here. The ground state energy (per site) of the Hamiltonian in Eq. (Topological materials with extensive flat-band surface states) is
$E_{0}(J,t)=-\frac{1}{N}\sum_{\vec{k}}\sqrt{A_{\vec{k}}^{2}+B_{\vec{k}}^{2}}$ (11)
where $N$ is the total number of sites. This quantity shows non-analytic behaviour as a function of $J$ and $t$ along some lines in the $J-t$ plane. These lines mark the boundaries of different phases (Fig. 1). The singularity for $t=1$ is presented in Fig. 2. There are six such phases, each characterised by specific topologies of nodal lines. The phase boundaries are also characterised by different types of non-analyticity in the ground state energy. Similar transitions between TNLSM phases have been observed earlier [28].
Figure 2: Derivatives of ground state energy calculated from Eq. (11) for $t=1$ and $N=300^{3}$.
Surface States: Let the eigenstates of the slab-Hamiltonian Eq. (4, 17) be $\psi_{n}(k_{x},k_{y})$ with eigenvalues $E_{n}$ where ${n}$ runs over $-L_{z},-L_{z}+1,\cdots L_{z}$. We identify numerically, those points $(k_{x},k_{y})$ for which there exists at least one ${n}$ with $E_{n}=0$ and with large values of $|\psi_{n}(k_{x},k_{y})|^{2}$ at ±$\pm L_{z}$ (Fig. 3). Fig. 1 describes the different types of surface states observed numerically in the six phases.
Figure 3: The probability distribution of all the $E=0$ states vs $n_{z}$ in phase III(left) and phase V(right).
From an analytic viewpoint, one finds that the Hamiltonian in Eq. (17) shows surface states when the condition
$\Delta_{1}^{2}+\Delta_{2}^{2}<1$ (12)
is satisfied (see Appendix section for details).
To calculate the winding number, we write the Hamiltonian ${\mathcal{H}}_{\vec{k}}$ of Eq. (1) as
${\mathcal{H}}_{\vec{k}}=R_{x}\sigma_{z}+R_{y}\sigma_{x}$ (13)
where we introduce a vector $\vec{R}$ with components
$R_{x}=J-t\cos k_{x}-t\cos k_{y}-\cos k_{z}$
$R_{y}=t\sin k_{x}+t\sin k_{y}+\sin k_{z}$ (14)
In view of the equality
$(R_{x}-\Delta_{1})^{2}+(R_{y}-\Delta_{2})^{2}=1$ (15)
we can see that as $k_{z}$ moves from $-\pi$ to $\pi$, the tip of the vector $\vec{R}$ moves over a circle with radius 1 and centre $(\Delta_{1},\Delta_{2})$. Therefore, the winding number $W$ will be $1$ or 00 according as $\Delta_{1}^{2}+\Delta_{2}^{2}$ is smaller or larger than 1 respectively. That the condition for $W=1$ is the same as the condition Eq. (12) for the existence of the surface states, indicates validity of the bulk-boundary correspondence theorem [29]. The regions of $W=1$ are shown for different phases in Fig. 4.
The dispersion relations are presented in the Appendix section.
Figure 4: Regions (coloured) for which the winding number is 1 on the $k_{x}$-$k_{y}$ plane with fully periodic boundaries. Note that for $J<1$, these states form a multiply connected region, and for $J>1$ they form one or more simply connected region(s). Also, these regions are the same as the regions for zero energy surface states in Fig. Fig. 1a under periodic boundary condition in XY plane.
Discussions: We study a system of spinless fermions described by the Hamiltonian (Eq. 16)
${\mathcal{H}}=\sum_{\vec{k}}\left[\left(J-t\cos k_{x}-t\cos k_{y}-\cos k_{z}\right)\sigma_{z}+\left(t\sin k_{x}+t\sin k_{y}+\sin k_{z}\right)\sigma_{x}\right]$ (16)
and show that as a function of $t$ and $J$, there appear several phases having different types of topological properties of nodal lines and zero-energy surface states. The phase boundaries are also marked by different types of non-analyticities. The principal result of this work is that one of these phases (that exists for $10<J<$ 1, $20<t<(1-J)/$ 2) has flat band surface states at zero energy. We mention that our Hamiltonian can be realised by suitably tuning a slab of TNLSM or by stacking layers each of which are described by Kitaev model on honeycomb lattice.
It will be interesting to analyse the behaviour of conductivity for our Hamiltonian and to look for materials which may be represented by it.
Appendix
Dispersion Relation The dispersion relation of the effective surface Hamiltonian of the surface $(0,0,1)$ is given with respect to $k_{y}$ for $k_{x}=0$, $k_{x}=\pi/2$ and $k_{x}=\pi$ in different phases in Fig (5-9). The calculation of the dispersion is done for $100$ layers perpendicular to the $z$-direction. The flat bands that are present in different phases correspond to the surface states, which are localized at surfaces $n_{z}=0$ and $n_{z}=100$. In the figures (5-9), we have shown only 40 bands around $E=0$. In Fig (5), we can see that flat band surface states are present for every $k_{x}$ and $k_{y}$ in phase I. [30, 31]
(a) $k_{x}=0$
(b) $k_{x}=\pi/2$
(c) $k_{x}=\pi$
Figure 5: Dispersion relation for Phase I : The dispersion relations are calculated for $J=0.2$ and $t=0.2$
(a) $k_{x}=0$
(b) $k_{x}=\pi/2$
(c) $k_{x}=\pi$
Figure 6: Dispersion relation for Phase II : The dispersion relations are calculated for $J=0.5$ and $t=0.5$
(a) $k_{x}=0$
(b) $k_{x}=\pi/2$
(c) $k_{x}=\pi$
Figure 7: Dispersion relation for Phase III : The dispersion relations are calculated for $J=0.5$ and $t=1.0$
(a) $k_{x}=0$
(b) $k_{x}=\pi/2$
(c) $k_{x}=\pi$
Figure 8: Dispersion relation for Phase IV : The dispersion relations are calculated for $J=2.0$ and $t=2.0$
(a) $k_{x}=0$
(b) $k_{x}=\pi/2$
(c) $k_{x}=\pi$
Figure 9: Dispersion relation for Phase V : The dispersion relations are calculated for $J=2.5$ and $t=1.0$
Surface States: We shall derive here the condition under which the Hamiltonian (see Eq. 7 of main text)
$H=\left(\begin{array}[]{cccccc}S&\sigma^{-}&0&0&\cdots&0\\ \sigma^{+}&S&\sigma^{-}&0&\cdots&0\\ 0&\sigma^{+}&S&\sigma^{-}&\cdots&0\\ \cdots&&&&&\\ 0&0&0&\cdots&\sigma^{+}&S\end{array}\right)$ (17)
admits of zero energy states at either surface $n_{z}\pm L$. Here, $S=\Delta_{1}\sigma_{z}+\Delta_{2}\sigma_{x}$ and $\sigma^{\pm}=-(\sigma_{z}\pm i\sigma_{x})/2$. This is an $N\times N$ matrix, each element of which is a $2\times 2$ matrix.
Since this matrix has the structure of Su-Schrieffer-Heeger Hamiltonian, we shall follow Asboth [29] and write the eigenvector of $H$ as a column matrix composed of the two-component column matrices $|x_{1}\rangle$, $|x_{2}\rangle$, $\cdots$ $|x_{N}\rangle$. The energy eigenvalue equations then become
$S|x_{1}\rangle+\sigma^{-}|x_{2}\rangle$ $=$ $E|x_{1}\rangle$ (18)
$\sigma^{+}|x_{k-1}\rangle+S|x_{k}\rangle+\sigma^{-}|x_{k+1}\rangle$ $=$ $E|x_{k}\rangle$
for  $\text{for}\;\;1<k<N$ (19)
$\sigma^{+}|x_{N-1}\rangle+S|x_{N}\rangle$ $=$ $E|x_{N}\rangle$ (20)
Surface states exist when these equations admits of solution with $E=0$ and either or both of $|x_{1}\rangle$ and $x_{N}\rangle$ non-zero. We note the identities
$\sigma^{+}S\sigma^{-}=0\;\;\;\;\;S^{2}=\lambda^{2}\underline{1},\mbox{with}\lambda=\sqrt{\Delta_{1}^{2}+\Delta_{2}^{2}}$
and put $E=0$ in Eqs (18,19) to get
[…]
This gives,
$\frac{1}{2}(1+\sigma_{y})|x_{N}\rangle=\lambda^{N-1}e^{i\phi(N-1)}|x_{1}\rangle$
where $\phi$ is defined by $\Delta_{1}+i\Delta_{2}=\lambda^{i\phi}$. When $\lambda<1$, one solution of this equation for large $N$ is $|x_{N}\rangle\approx 0$ even when $\langle x_{1}|x_{1}\rangle\sim 1$. Hence, Eq. (20) is satisfied, although approximately, with $E=0$, indicating the presence of zero-energy state at the edge $|x_{1}\rangle$.
Critical lines As mentioned in Fig. 1 of the main text, the non-analyticity at the phase boundaries is in the form of a discontinuity in $\partial^{3}E_{0}/\partial t^{3}$ in some cases and in the form of divergence of $\partial^{3}E_{0}/\partial t^{3}$ in the other cases. The amount of discontinuity varies along the boundary line. The line $J=1$ forms the boundary (II, V) and the one (III, IV). We show the variation of the amount of discontinuity along this line in Fig. (10(a)). Similarly, the line $J-2t=-1$ forms the boundary (II, III) and the one (IV, V). The variation of the amount of discontinuity along this line is shown in Fig. (10(b)). On the other hand, the third derivative $\partial^{3}E_{0}/\partial t^{3}$ diverges algebraically along the boundaries (I, II) and (V, VI), with the exponent depending upon the direction of approach. This is shown in Figs. (11, 12).
(a) $J=1$ critical line
(b) Phase boundary of IV and V/II and III
Figure 10: Amount of discontinuity vs $t$ : The behaviour of the amount of discontinuity in the third derivative of the ground state energy with the change of $t$ is shown in these plots. Note that the points lie on a smooth line apart from the point for $t=1$.
Figure 11: Algebraic divergence of $\partial^{3}E_{0}/\partial J^{3}$ at the critical line $J-2t=1$ : The dots are the numerical data and the green line is the fitted curve of the form $ax^{b}$. The critical exponent is $b=0.11$ when the critical line is approached from below and $b=0.14$ when the critical line is approached from above. $t$ is fixed at $0.2$. Both $x$ and $y$ axes are in log scale.
Figure 12: Algebraic divergence of $\partial^{3}E_{0}/\partial J^{3}$ at the critical line $J+2t=1$ : The dots are the numerical data and the green line is the fitted curve of the form $ax^{b}$. The critical exponent is $b=0.19$ when the critical line is approached from below and $b=0.22$ when the critical line is approached from above. $t$ is fixed at $0.3$. Both $x$ and $y$ axes are in log scale.
Dispersion relations of Phase I with disorder We find that a random fluctuation of the hopping amplitude around 1.0 by 10% does not destroy the zero energy surface states (Fig. 13).
(a) $k_{x}=0$
(b) $k_{x}=\pi/2$
(c) $k_{x}=\pi$
Figure 13: Dispersion relation for Phase I : The dispersion relations are calculated for $J=0.2$ and $t=0.25$. The disorder is put in the Z direction where it fluctuates uniformly and randomly between $0.9$ and $1.1$.